OpenStudy (anonymous):

An artist wants to make alabaster pyramids using a block of alabaster with a volume of 576 cubic inches. She plans to make each pyramid with a square base area of 3 square inches and a height of 4 inches. At most, how many pyramids can the artist make from the block of alabaster?

OpenStudy (anonymous):

A. 136 B. 144 C. 154 D. 166 E. 172

OpenStudy (anonymous):

@AnswerMyQuestions @leahhhmorgannn

OpenStudy (leahhhmorgannn):

I don't do math, sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

im thinking 144

OpenStudy (anonymous):

\(\sf V = b^2 \times \dfrac{1}{3} h\) Plug in what we know: \(\sf V = 3^2 \times \dfrac{1}{3} (4)\) Simplify

OpenStudy (anonymous):

so 12?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Now divide that to 576.

OpenStudy (anonymous):

\(\sf 576 \div 12=~?\)

OpenStudy (anonymous):

48?

OpenStudy (anonymous):

Yep, that's our answer.

OpenStudy (anonymous):

but thats not a choice